OpenStudy (anonymous):

I need help please on how to use sigma notations to write the sum[5(1/8)+3]+[5(2/8)+3]+...+[5(8/8)+3]

OpenStudy (anonymous):

\[5\sum_{i = 1}^{8}\left( \frac{ i }{ 8 } + 3 \right)\]

OpenStudy (anonymous):

i can answer any questions you have about this

OpenStudy (anonymous):

Thanks but what are the step to get to that can you explain

OpenStudy (anonymous):

i just looked at it to deduce it. no formulas. do you know what the numbers above and below the sigma mean?

OpenStudy (anonymous):

actually it's wrong

OpenStudy (anonymous):

tried to be smart by taking out the 5 sine it multiplies every i/8 term but i forgot about the +3. this is the real answer:\[\sum_{i =1 }^{8}\left( 5\frac{ i }{ 8 } + 3\right)\]

OpenStudy (anonymous):

since*

OpenStudy (anonymous):

Ah okay thanks I see how you did it now thank you

OpenStudy (anonymous):

glad to help ^_^

OpenStudy (anonymous):

Could you help me on how to solve this problem

OpenStudy (anonymous):

(20 sigma notation I=1) 2i